OpenStudy (uri):

Stats help: A card is drawn from a well shuffled pack of cards.Then a second card is drawn.Find the probability that both are spades if first card is not replaced before the second is drawn.

OpenStudy (souvik):

are there jokers?

OpenStudy (uri):

That's the question.After that it says nothing.

OpenStudy (souvik):

lets take there is no joker.. let A is the probability of first card being a spade... P(A)=13C1/52C1 agree?

OpenStudy (uri):

Okay so there are like 52 cards in one set and from them 13 are spade?

OpenStudy (souvik):

yes...!

OpenStudy (uri):

both are spade it says..?

OpenStudy (souvik):

let B be the probability of second card being a spade... so P(B)=12C1/51C1 as there is no replacement agree?

OpenStudy (uri):

12? where is it coming from? spade is 13 no?

OpenStudy (souvik):

first card is not replaced ...so there is 12 other spades.

OpenStudy (uri):

I don't get it -_-

hartnn (hartnn):

ok, 13 spades in 52 cards when 1st card is drawn, P(spade) = # of spades / total # of cards = 13/52 clear till here ?

OpenStudy (uri):

Yes :3

hartnn (hartnn):

good! now we have drawn out one of the spade out of 13 spades, how many remain ?

OpenStudy (uri):

12.

hartnn (hartnn):

right, and in total number, 51 will remain. so, P (2nd card to be spade) = # of remaining spades / # of remaining total cards = 12/51 clear ?

OpenStudy (uri):

Yes clear.

hartnn (hartnn):

now last part, probability that both are spades = P (1st card is spade ) AND P (2nd card is spade) since there's a AND , we will multiply these events (probability of events) probability that both are spades = (13/52) times (12/51) = ... ?

OpenStudy (uri):

0.058823

OpenStudy (souvik):

correct !

hartnn (hartnn):

or you can keep it as 3/51

OpenStudy (uri):

Oh that's all :3,thankyou guys :3

OpenStudy (souvik):

you are welcome...

hartnn (hartnn):

welcome ^_^